angles 1 and 2 are supplementary . Angle 2 is 30 degrees more than twice the measure of Angle 1
Answer by nerdybill(5404)

angles 1 and 2 are supplementary . Angle 2 is 30 degrees more than twice the measure of Angle 1
.
If two angles are supplementary, then the sum of the two angles is 180 deg.
.
Let x = measure of angle 1
then
2x+30 = measure of angle 2
.
x + 2x+30 = 180
3x+30 = 180
3x = 150
x = 50 degrees (angle 1)
.
angle 2:
2x+30 = 2(50)+30 = 100+30 = 130 deg (angle 2)